COM 1205 Software Design and Development — Syllabus Fall 1996 Professor Karl. J. Lieberherr

The syllabus is subject to change based on class reaction and projects selected by the students. The AP book is the text book: Adaptive Object-Oriented Software: The Demeter Method.

This course provides state-of-the-art techniques and concepts for object-oriented software design and development. You learn the Demeter Method for object-oriented software development which will hopefully make you at least 4 times as productive as a Java software developer compared to software development without the method. In the first few weeks you learn the method which you apply in the remaining weeks to your project. The project will be done in groups.

Assume that you could do your project in 200 hours of Java programming without using a good method. Instead of working hard, we will work smart. We will spend 50 hours on learning the Demeter Method and then you spend 50 hours on doing your project in Java using the Demeter Method. Since you will be at least 4 times as productive, you will still finish your 200 hour project in 50 hours only. Instead of spending 200 hours on the course, which would be excessive, you can accomplish the same in a total of 100 hours AND learn very useful object-oriented technology.

Designing and programming will be done in a structure-shy, grammar-based, object-oriented style. All assignments and the project will be written directly or indirectly in Java. If you choose, you may use an experimental tool, called Demeter/Java.

This course does not have Java as a prerequisite, however you are expected to learn a small subset of Java from the recommended text book or from another good Java book of your choice.

1. Week: Introduction to adaptive software development. Pattern: Structure-shy Traversal.

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/AOP/Structure-Shy-Traversal

Pattern: Context.

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/AOP/Context

Class dictionaries and Unified Modeling Notation.

AP book chapters: selections from 1-4.

2. Week: Pattern: Class Graph.

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/AOP/ClassGraph

Writing simple adaptive programs using traversals and visitors. Class dictionary design.

AP book chapters: selections from 5-8.

3. Week: Pattern: Structure-shy Object.

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/AOP/Structure-Shy-Object Propagation directives. Decoupling classes: Law of Demeter Relationship to adaptive software. Notations for documenting class libraries and frameworks. The class dictionary notation (graphical and textual). Design rule checking of class dictionaries. Class dictionaries as customizers for adaptive programs.

AP book chapters: selections from 8-11.

4. Week: Pattern: Inventor’s paradox

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/AOP/IP.html

More complex uses of visitors: multiple visitors and nested visitors. Garbage collection. Improving the reusability of software designs. Parameterized class definitions. Style rules for class dictionaries.

AP book chapters: selections from 10-12.

5. Week: The Demeter Method for AP. Design notations for behavior. Developing a growth plan for implementation and testing. Design with maintenance in mind.

AP book chapters: 13.

6. Week: Midterm. Design notations for behavior. Exception handling design. Testing of object-oriented software.

7. Week: Graphical user-interface development with the Java framework AWT (Abstract Window Toolkit).

8. Week: Graphical user-interface development with the Java framework AWT (continued). The Unified Modeling Language (continued).

9. Week: Further design patterns for adaptive object-oriented software development.

http://www.ccs.neu.edu/research/demeter/adaptive-patterns/

10. Week: A comparison of object-oriented software development methodologies.

11. Exam week: Final.